OpenStudy (anonymous):

HELP PLEASE A segment with endpoints A (2, 6) and C (5, 9) is partitioned by a point B such that AB and BC form a 3:1 ratio. Find B. (2.33, 6.33) (3.5, 10.5) (3.66, 7.66) (4.25, 8.25)

OpenStudy (anonymous):

Okay. the interval of the x parts is 2 to 5, giving us a distance of 3. 3/4 = 1 part so 3*3/4 = 3 parts, agreed?

OpenStudy (anonymous):

this is how .... you work the x and y parts or Bx = 5 - 1*3/4 either way gets us to the same place eh

OpenStudy (anonymous):

or in this case it seems like we only need to compare the x parts to match an option

OpenStudy (anonymous):

So, it would be (4.25, 8.25 )

OpenStudy (anonymous):

Hope this helped! Have a great day! @jaja420

OpenStudy (anonymous):

THANKS!!!!!